Can you please explain...instructions.
Tickets for a play at the community theater cost $20 for an adult and $8 for a child...if 220 tickets were sold and the total receipts were $3200, how many of each type of ticket were sold.
20x + 8y = 3200

Answer by ptaylor(1658)

OK You can work this problem using two unknowns:
Let x=number of adult tickets sold
And y=number of child tickets sold
Then as you say:
20x+8y=3200--------------Equation 1.
The other equation is:
x+y=220-----------------Equation 2.
(1) 20x+8y=3200
(2) x+y=220
multiply equation (2) by 8
(1) 20x+8y=3200
(2a) 8x+8y=1760
subtract (2) from (1)
12x=1440 divide both sides by 12
x=120-----------------------number of adult tickets sold
Now substitute x=120 into (2)
120+y=220 subtract 120 from both sides
y=100-----------------------number of child tickets sold
CK
120*20+100*8=3200
2400+800=3200
3200=3200
You can also work this problem using one unknown:
Let x=number of adult tickets sold
Then 220-x=number of child tickets sold
Now we are told that cost of adult tickets sold plus cost of child tickets sold equals $3200. So our equation to solve is:
20x+8(220-x)=3200 get rid of parens
20x+1760-8x=3200 subtract 1760 from both sides
12x=1440 divide both sides by 12
x=120--------------------------------number of adult tickets sold
22-x=220-120=100----------------------number of child tickets sold


Hope this helps----ptaylor